An astronaut's pack weighs 10.8 N when she is on earth but only 4.15 N when she is at the surface of moon.

Part A

What is the acceleration due to gravity on this moon?

Part B

What is the mass of the pack on this moon?
Answer #1

here,

weight on earth , we = 10.8 N

and weight at the moon surface , Wm = 4.15 N

A)

let the accelration due to gravity on moon be g'

g'/g = mg'/mg= wm /we

g'/g = 4.15 /10.8

g' = 9.8 * 4.15 /10.8 = 3.77 m/s^2

the accelration due to gravity on moon is 3.77 m/s^2

B)

the mass of the pack , m =
wm/g' = 4.15/3.77 = 1.1 kg

the mass of the pack on moon is 1.1 kg